OpenStudy (anonymous):

the square of one number is 2 less than twice the square of another. the sum of the squares of the two numbers is 25. find the numbers. help please.

OpenStudy (mathstudent55):

Choose two variables to represent the two numbers.

OpenStudy (anonymous):

can i use x and y?

OpenStudy (mathstudent55):

Of course.

OpenStudy (mathstudent55):

Now let the first number be x. How do you express "the square of one number" using x?

OpenStudy (mathstudent55):

the square of one number is 2 less than twice the square of another \(x^2\) = \(2y^2 - 2\)

OpenStudy (mathstudent55):

From the first statement, we have the equation: \(x^2 = 2y^2 - 2\) The second sentence is: "the sum of the squares of the two numbers is 25." Now we need to translate the second sentence into an equation. Can you try doing it?

OpenStudy (anonymous):

\[x ^{2}+2y ^{2}=25?\]

OpenStudy (mathstudent55):

No, close, but there is no 2. Just \(x^2 + y^2 = 25\)

OpenStudy (mathstudent55):

Now we have two equations that we need to solve as a system of equations: \(x^2 = 2y^2 - 2\) \(x^2 + y^2 = 25\) Let's solve the second equation for \(x^2\). \(x^2 = 25 - y^2\) Ok so far?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Since we know \(x^2 = 2y^2 - 2\), and \(x^2 = 25 - y^2\) then that means \(2y^2 - 2 = 25 - y^2\) Add 2 to both sides, and add \(y^2\) to both sides. \(3y^2 = 27\) Divide both sides by 3 \(y^2 = 9\)

OpenStudy (anonymous):

ok i understand that

OpenStudy (mathstudent55):

Now we can find y.

OpenStudy (mathstudent55):

In general, if you have an equation of the form \(x^2 = k\), where k is a number, then the solution is \(x = \sqrt{k}\) or \(x = -\sqrt{k} \)

OpenStudy (mathstudent55):

In this case, we have \(y^2 = 9\), so we get \(y = \sqrt{9}\) or \(y = -\sqrt{9} \) which gives us \( y = 3\) or \(y = -3\)

OpenStudy (mathstudent55):

Now that we have a solution for y, we plug the values of y into one of our equations to get x. We need to do it for both values of y. Let's us the second equation: \(x^2 + y^2 = 25\)

OpenStudy (mathstudent55):

Let's start with y = 3: \(x^2 + 3^2 = 25\) \(x^2 + 9 = 25\) \(x^2 = 16\) \(x = 4\) or \(x = -4\) This means that for y = 3, we have two solutions to the system of equations: x = 4, y = 3 and x = -4, y = 3

OpenStudy (mathstudent55):

Now we do the same for y = -3: \(x^2 + (-3)^2 = 25\) \(x^2 + 9 = 25\) \(x^2 = 16\) \(x = 4\) or \(x = -4\) This means that for y = -3, we have two solutions to the system of equations: x = 4, y = -3 and x = -4, y = -3

OpenStudy (mathstudent55):

There are 4 pairs of numbers that work, so there are 4 answers to the question: 3, 4 3, -4 -3, 4 -3, -4

OpenStudy (anonymous):

will there always be a positive and negative solution?

OpenStudy (mathstudent55):

Not always, but when you deal with squares, sometimes you do get them.

OpenStudy (anonymous):

thank you!

OpenStudy (mathstudent55):

wlcm